OpenStudy (brydenwright):

The given measurements may or may not determine a triangle. If not, then state that no triangle is formed. If a triangle is formed, then use the Law of Sines to solve the triangle, if it is possible, or state that the Law of Sines cannot be used. B = 126°, c = 7, b = 12

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Do you know the Law of Sines and how to use it?

OpenStudy (evoker):

I think the 12 is opposite the 126 right?

OpenStudy (brydenwright):

I know the Law of Sines, but I don't know how to figure out if a triangle can be made or not.

OpenStudy (mathstudent55):

You are correct. My mistake. Here is the corrected drawing. 

OpenStudy (mathstudent55):

First let's see if you can use the Law of Sines.

OpenStudy (mathstudent55):

To use the Law of Sines, you need to have a given side and its opposite angle. That will establish the ratio of the Law of Sines.

OpenStudy (mathstudent55):

We do have angle B and side b. \(\dfrac{b}{\sin B}= \dfrac{c}{\sin C} \)

OpenStudy (mathstudent55):

We know b, c, and B. We can solve for C.

OpenStudy (mathstudent55):

Find C. Then find A using the sum of angles. Then find side "a" using the Law of Sines again.